OpenStudy (anonymous):

ok i got my answer to be 75/3.... integral 1 to 2 (1+2y)^2dy

OpenStudy (anonymous):

It should be 49/3

OpenStudy (anonymous):

not sure where i went wrong

OpenStudy (anonymous):

integral 4y^2 + 4y + 1 4y^3/3 +4y^2/2 + y plug in 1 and 2

OpenStudy (anonymous):

heres what i did/.....\[\int\limits_{1}^{2} (1+2y)^{2} dy\] \[(1+2y)(1+2y)= 1+4y=4y ^{2}\]

OpenStudy (anonymous):

1y+(4y^2/2)+(4y^3/3) with integral 1 to 2

OpenStudy (anonymous):

i got the first part to 13/3+62/3...

OpenStudy (anonymous):

i meant minus!!!!ugghhh now i see

OpenStudy (anonymous):

[32/3+8+2]-[4/3+2+1]

OpenStudy (anonymous):

ok now i know thanks i see where i got mess up